# Galactic Archaeology with Gaia

### v1

@@ 3.1 Selecting GS/E members
Perhaps the most important first step in such studies is to understand how exactly we select the GS/E stars, as any deduction of its properties (e.g. total mass, metallicity etc) will be significantly impacted by any selection effects. On the one hand, the proximity of this structure to the solar neighborhood means we have significant numbers of potential GS/E stars to study. However, this also means that the GS/E is located in a region close to the stellar disc, and is thus prone to contamination. In the Belokurov et al. (2018b) discovery paper the authors use a GMM method to simultaneously model the different Milky Way components in velocity space, but even this advance over simple kinematic cuts cannot guarantee a ‘pure’ sample of GS/E stars. Unsurprisingly, more recent works have looked to action-angle space to locate the GS/E stars (e.g. Feuillet et al., 2020; Lane et al., 2022), which should result in a cleaner selection than velocity information alone. Other works have made use of the exquisite APOGEE data to perform a purely chemical selection of GS/E stars, whereby a cut in the plane of [Mg/Mn] versus [Al/Fe] is assumed to cleanly separate accreted and in-situ stars (e.g. Das et al., 2020; Carrillo et al., 2022; Lane et al., 2023).
@@ 3.2 Halo kinematics and the orbital make-up of the GS/E debris
Similarly, Lancaster et al. (2019) rely on an identical two-component mixture to model a sample of previously identified BHB stars reaching far beyond the Solar neighborhood. Again, one model component is allowed to have a double-peaked $v_{r}$ distribution with an unconstrained separation between the peaks. They demonstrate that a single-Gaussian description of the velocity distribution of the Galactic stellar halo is a gross over-simplification. Lancaster et al. (2019) show that out to $r\approx 25$ kpc from the Galactic centre, such a model does not fit the data correctly. They show that within this region, the BHB kinematics are best described by an equal contribution from a roughly isotropic halo and a radially anisotropic component, which they attribute to the GS/E merger. They also detect a decline in the separation between the radial lobes with Galacto-centric distance $r$, as expected from merger simulations (see Fig. 2). Lancaster et al. (2019) determine that beyond $r\approx 25$ kpc, the fractional contribution of the bi-modal $v_{r}$ component is consistent with zero. These findings are in agreement with the detection (see e.g. Sesar et al., 2011; Watkins et al., 2009; Deason et al., 2011b) and the interpretation (see e.g. Deason et al., 2013a, 2018) of the stellar halo density break at a similar distance.
The appeal of BHBs is that despite being quite rare, they are a robust standard candle that can be picked out of a mix of stellar populations cleanly and easily. Compared to BHBs, K-giants are a more numerous and brighter halo tracer, at the expense of being a less accurate distance indicator. Bird et al. (2019) and Bird et al. (2021) use a large number of K-giants to probe the global anisotropy profile of the Galactic stellar halo from 5 to 100 kpc. Their single multivariate Gaussian model shows qualitatively similar behaviour to the singe-component model of Lancaster et al. (2019). However, they report a higher orbital anisotropy for K-giants compared to the BHBs. Moreover, unlike the anisotropy of the BHBs which shows a pronounced decline around $r\approx 20$ kpc as discussed above, the K-giants’ $\beta$ remains more or less flat across the whole range of Galactocentric distances probed.
@@
Figure 12: Top row: Strength of correlation between the longitudinal and latitudinal components of proper motion in Galactic coordinates for a model with a strong radial anisotropy (left) and Gaia DR2 RR Lyrae (right). Adapted from Iorio and Belokurov (2019). Bottom row: Fractional contribution of the radially-biased halo component in the Gaia RR Lyrae as a function of the Galactocentric distance (left) and the radial velocity offset (the separation between the two Gaussians is twice this value) in the kinematic model of radially-biased halo component of the Gaia RR Lyrae (right). [Adapted from Iorio and Belokurov (2021)].
@@
A minor tweak was introduced by Iorio and Belokurov (2019) to their earlier model (Iorio et al., 2018) in order to tilt the halo slightly out of the Milky Way’s disc plane (see also Han et al., 2022a). This non-axisymmetry is warranted by the RR Lyrae number counts and is likely related, at least in part, to the two previously identified large-scale sub-structures, the Virgo Overdensity (VOD Jurić et al., 2008; Vivas et al., 2016) and the Hercules Aquila Cloud (HAC Belokurov et al., 2007; Simion et al., 2014). Curiously, VOD and HAC are shown to be closely related by Simion et al. (2019) who, based on orbital analysis, argue that both originated in the GS/E merger. A similar conclusion was reached by Balbinot and Helmi (2021). The tilt in the GS/E debris cloud with respect to the disc as exemplified by the HAC-VOD asymmetry is examined by Han et al. (2022b) who point out that such an asymmetry can not survive for very long in an axi-symmetric potential of the Galaxy. They argue that to maintain the stellar halo’s tilt over a long period of time, a similarly tilted dark matter component is required. This idea is fully fleshed out in Han et al. (2023a) and Han et al. (2023b) where the misaligned dark matter halos are analysed and shown to be ubiquitous in Cosmological simulations. Such tilted dark matter halo components are damaging to the Galactic discs, causing them to warp and flare. In a similar vein, Davies et al. (2023) argue that the evolution of the GS/E debris in an axi-symmetric potential would lead to discernible striation in the integrals-of-motion space once the stellar halo is tugged by the passing LMC. The fact that no such characteristic substructure is observed (although see Belokurov et al., 2023, for the detection and interpretation of high-frequency substructure in the GS/E debris) could be interpreted as evidence for a non-axisymmetric component in the Milky Way’s dark matter distribution. From the analysis in Dillamore et al. (2022), it is clear that tilts such as those proposed in Han et al. (2023a) and Davies et al. (2023) are common in galaxies recovering from a significant merger.
@@
Iorio and Belokurov (2019) also uncover a tight kinematic bond amongst most of the inner halo RR Lyrae. They show that within $r\approx 25$ kpc from the Galactic centre, the stellar motions are correlated: the kinematics of the bulk of the RR Lyrae is consistent with a nearly radial motion, one of the key characteristics of the GS/E. Iorio and Belokurov (2021) move beyond demonstration to modeling of the kinematic properties of the Gaia RR Lyrae sample (see Fig. 12). Their model contains three distinct components: a rotating disc-like population, a quasi-isotropic one, and finally, a strongly radially anisotropic population corresponding to the GS/E. The modeling approach is inspired by an earlier work of Wegg et al. (2019) where 5D data with missing line-of-sight velocity is modeled under the assumption of axi-symmetry to extract the properties of the RR Lyrae velocity ellipsoid. For the radially anisotropic GS/E population, Iorio and Belokurov (2021) rely on the functional form introduced in Necib et al. (2019) and Lancaster et al. (2019) and map out the separation between the GS/E radial velocity bumps inside $r\approx 30$ kpc. The separation reaches $\approx 500$ km/s around $r\approx 5$ kpc and drops to zero around $r\approx 25$ kpc. The GS/E contribution to halo RR Lyrae is maximal around the Sun at $\sim 75\%$ but decreases to $\sim 20\%$ inside $r\approx 5$ kpc and outside $r\approx 20$ kpc. Curiously, the properties of the isotropic component are not uniform throughout: Iorio and Belokurov (2019) show that the central $\approx 10$ kpc are dominated by a more metal-rich RR Lyrae population with a higher fraction of rare High Amplitude Short Period pulsators (see Fiorentino et al., 2015; Belokurov et al., 2018a). Given its properties, this centrally concentrated RR Lyrae population is likely of in-situ origin.
@@ 3.3 Chemical trends
Even though the GS/E progenitor is no longer intact, its pre-disruption properties can be studied now, provided the constituent stars are identified (see Sections 3.1 and  3.2). As the GS/E stars flood the Solar neighborhood, as aptly pointed out by Hasselquist et al. (2021), the community had been exploring the GS/E’s chemical trends long before we had realised these stars were all of common origin (see Nissen and Schuster, 2010; Hawkins et al., 2015; Fernández-Alvar et al., 2018; Hayes et al., 2018). Thus, the main trends, i.e. the downward slope of the $\alpha$-[Fe/H] track (Nissen and Schuster, 2010) and the constant low [Al/Fe] ratio (Hawkins et al., 2015) were known before the Gaia data started to arrive. These trends are updated, improved upon, and contrasted with those in surviving dwarf galaxies in Hasselquist et al. (2021) (see Fig. 13). They show that compared to the four most massive current satellites of the Milky Way, namely the LMC, the SMC, Sgr (holding on for dear life), and Fornax, the GS/E progenitor attains demonstrably higher [Mg/Fe] at $-1.5<$[Fe/H]$<-1$ (but still some 0.1 dex below the disc sequence at the same metallicity). Clearly, the GS/E dwarf was more efficient at converting gas into stars compared to the present-day survivors. Hasselquist et al. (2021) hypothesize that this is due to its proximity to the Milky Way. Equally, the environment may have also played a crucial role in shaping the SFHs of the Magellanic Clouds which appear to have lived a mostly dull and lethargic life in the suburbs of the Galaxy. However, having formed next to a bigger galaxy did not end well for the GS/E progenitor: its life was cut short, although none of the models in the literature reveal signs of abrupt truncation in star formation (possibly due to the rigidity of the SFH approximation).
@@
[…]
@@
Most of the chemical studies above use the exquisite APOGEE data. However, APOGEE is somewhat limited in its ability to measure heavy element abundances, and, in particular, has only scarce data on neutron-capture tracers (only Ce is listed amongst the relatively reliably measured species as part of the APOGEE DR17). The information on $s$- and $r$-process had to come from elsewhere, namely from the data releases of the GALAH survey (De Silva et al., 2015; Buder et al., 2018, 2021) and individual follow-up studies. These challenging measurements (due to a small number of weak lines) yielded a result not many had predicted: a noticeable enhancement in the $r$-process abundances in the GS/E stars compared to the rest of the halo (including the in-situ stars where available at comparable metallicities) and most of the surviving dwarf satellites (Aguado et al., 2021; Matsuno et al., 2021). Measuring the rather low [Ba/Eu]$\approx 0$ in their follow-up VLT spectra, Aguado et al. (2021) deduce that the neutron-capture production in the GS/E progenitor proceeded mostly via the $r$-process channel. Note that while [Eu/Fe]$\approx 0.65$ and [Eu/Mg]$\approx 0.3$ are at the upper end of what is observed, these values are still noticeably below the extreme $r$-process enhancement detected in one ultra-faint dwarf, namely Reticulum 2 (Ji et al., 2016) which reaches [Eu/Fe]$>1$, but similar to levels in other $r$-process enhanced UFDs such as Tucana 3 (Hansen et al., 2017) and Grus 2 (Hansen et al., 2020). Focusing on the tracks in [Eu/Mg]-[Mg/Fe] space for the GS/E stars with [Fe/H]$>-1.5$ in GALAH, Matsuno et al. (2021) conclude that the bulk of $r$-process in the dwarf must be due to neutron star mergers. Naidu et al. (2022) observe GS/E stars together with targets they designate “Kraken” with Magellan and measure (amongst other lines) Eu, Ba, and Mg abundances. At similar [Mg/H], the GS/E has 0.3 dex higher [Eu/Mg]. Naidu et al. (2022) assume that the “Kraken” had a more rapid star formation activity and conclude that the enrichment due to the neutron star mergers must be delayed by some 0.5 Gyr. Note that most “Kraken” stars analyzed in this study would also pass the cuts for the Milky Way pre-disc in-situ population (Aurora — see 4.3.1). Replacing “Kraken” with “early Milky Way”, the arguments presented in Naidu et al. (2022) remain valid, or perhaps, become even more coherent.

### v2

@@ 3.1 Selecting GS/E members
Perhaps the most important first step in such studies is to understand how exactly we select the GS/E stars, as any deduction of its properties (e.g. total mass, metallicity etc) will be significantly impacted by any selection effects. On the one hand, the proximity of this structure to the solar neighourhood means we have significant numbers of potential GS/E stars to study. However, this also means that the GS/E is located in a region close to the stellar disc, and is thus prone to contamination. In the Belokurov et al. (2018b) discovery paper the authors use a GMM method to simultaneously model the different Milky Way components in velocity space, but even this advance over simple kinematic cuts cannot guarantee a ‘pure’ sample of GS/E stars. Unsurprisingly, more recent works have looked to action-angle space to locate the GS/E stars (e.g. Feuillet et al., 2020; Lane et al., 2022), which should result in a cleaner selection than velocity information alone. Other works have made use of the exquisite APOGEE data to perform a purely chemical selection of GS/E stars, whereby a cut in the plane of [Mg/Mn] versus [Al/Fe] is assumed to cleanly separate accreted and in-situ stars (e.g. Das et al., 2020; Carrillo et al., 2022; Lane et al., 2023).
@@ 3.2 Halo kinematics and the orbital make-up of the GS/E debris
Similarly, Lancaster et al. (2019) rely on an identical two-component mixture to model a sample of previously identified BHB stars reaching far beyond the Solar neighourhood. Again, one model component is allowed to have a double-peaked $v_{r}$ distribution with an unconstrained separation between the peaks. They demonstrate that a single-Gaussian description of the velocity distribution of the Galactic stellar halo is a gross over-simplification. Lancaster et al. (2019) show that out to $r\approx 25$ kpc from the Galactic centre, such a model does not fit the data correctly. They show that within this region, the BHB kinematics are best described by an equal contribution from a roughly isotropic halo and a radially anisotropic component, which they attribute to the GS/E merger. They also detect a decline in the separation between the radial lobes with Galacto-centric distance $r$, as expected from merger simulations (see Fig. 2). Lancaster et al. (2019) determine that beyond $r\approx 25$ kpc, the fractional contribution of the bi-modal $v_{r}$ component is consistent with zero. These findings are in agreement with the detection (see e.g. Sesar et al., 2011; Watkins et al., 2009; Deason et al., 2011b) and the interpretation (see e.g. Deason et al., 2013a, 2018) of the stellar halo density break at a similar distance.
The appeal of BHBs is that despite being quite rare, they are a robust standard candle that can be picked out of a mix of stellar populations cleanly and easily. Compared to BHBs, K-giants are a more numerous and brighter halo tracer, at the expense of being a less accurate distance indicator. Bird et al. (2019) and Bird et al. (2021) use a large number of K-giants to probe the global anisotropy profile of the Galactic stellar halo from 5 to 100 kpc. Their single multivariate Gaussian model shows qualitatively similar behaviour to the single-component model of Lancaster et al. (2019). However, they report a higher orbital anisotropy for K-giants compared to the BHBs. Moreover, unlike the anisotropy of the BHBs which shows a pronounced decline around $r\approx 20$ kpc as discussed above, the K-giants’ $\beta$ remains more or less flat across the whole range of Galactocentric distances probed.
@@
Figure 12: Top row: Strength of correlation between the longitudinal and latitudinal components of proper motion in Galactic coordinates for a model with a strong radial anisotropy (left) and Gaia DR2 RR Lyrae (right) (Iorio and Belokurov, 2019). Bottom row: Fractional contribution of the radially-biased halo component in the Gaia RR Lyrae as a function of the Galactocentric distance (left) and the radial velocity offset (the separation between the two Gaussians is twice this value) in the kinematic model of radially-biased halo component of the Gaia RR Lyrae (right) (Iorio and Belokurov, 2021). [Adapted from Iorio and Belokurov (2019, 2021)].
@@
A minor tweak was introduced by Iorio and Belokurov (2019) to their earlier model (Iorio et al., 2018) in order to tilt the halo slightly out of the Milky Way’s disc plane (see also Han et al., 2022a). This non-axisymmetry is warranted by the RR Lyrae number counts and is likely related, at least in part, to the two previously identified large-scale sub-structures, the Virgo Overdensity (VOD Jurić et al., 2008; Vivas et al., 2016) and the Hercules Aquila Cloud (HAC Belokurov et al., 2007; Simion et al., 2014). Curiously, VOD and HAC are shown to be closely related by Simion et al. (2019) who, based on orbital analysis, argue that both originated in the GS/E merger. A similar conclusion was reached by Balbinot and Helmi (2021). Moreover, Perottoni et al. (2022) show that, in addition to their dynamical link, the VOD and HAC are chemically indistinguishable from the prototypical GS/E population. The tilt in the GS/E debris cloud with respect to the disc as exemplified by the HAC-VOD asymmetry is examined by Han et al. (2022b) who point out that such an asymmetry cannot survive for a very long in an axi-symmetric potential of the Galaxy. They argue that to maintain the stellar halo’s tilt over a long period of time, a similarly tilted dark matter component is required. This idea is fully fleshed out in Han et al. (2023a) and Han et al. (2023b) where the misaligned dark matter halos are analysed and shown to be ubiquitous in Cosmological simulations. Such tilted dark matter halo components are damaging to the Galactic discs, causing them to warp and flare. In a similar vein, Davies et al. (2023) argue that the evolution of the GS/E debris in an axi-symmetric potential would lead to discernible striation in the integrals-of-motion space once the stellar halo is tugged by the passing LMC. The fact that no such characteristic substructure is observed (although see Belokurov et al., 2023, for the detection and interpretation of high-frequency substructure in the GS/E debris) could be interpreted as evidence for a non-axisymmetric component in the Milky Way’s dark matter distribution. From the analysis in Dillamore et al. (2022), it is clear that tilts such as those proposed in Han et al. (2023a) and Davies et al. (2023) are common in galaxies recovering from a significant merger.
@@
Iorio and Belokurov (2019) also uncover a tight kinematic bond amongst most of the inner halo RR Lyrae. They show that within $r\approx 25$ kpc from the Galactic centre, the stellar motions are correlated: the kinematics of the bulk of the RR Lyrae is consistent with a nearly radial motion, one of the key characteristics of the GS/E (see also Ablimit et al. 2022). Iorio and Belokurov (2021) move beyond demonstration to modelling of the kinematic properties of the Gaia RR Lyrae sample (see Fig. 12). Their model contains three distinct components: a rotating disc-like population, a quasi-isotropic one, and finally, a strongly radially anisotropic population corresponding to the GS/E. The modelling approach is inspired by an earlier work of Wegg et al. (2019) where 5D data with missing line-of-sight velocity is modeled under the assumption of axi-symmetry to extract the properties of the RR Lyrae velocity ellipsoid. For the radially anisotropic GS/E population, Iorio and Belokurov (2021) rely on the functional form introduced in Necib et al. (2019) and Lancaster et al. (2019) and map out the separation between the GS/E radial velocity bumps inside $r\approx 30$ kpc. The separation reaches $\approx 500$ km/s around $r\approx 5$ kpc and drops to zero around $r\approx 25$ kpc. The GS/E contribution to the halo RR Lyrae is maximal around the Sun at $\sim 75\%$ but decreases to $\sim 20\%$ inside $r\approx 5$ kpc and outside $r\approx 20$ kpc. Curiously, the properties of the isotropic component are not uniform throughout: Iorio and Belokurov (2019) show that the central $\approx 10$ kpc are dominated by a more metal-rich RR Lyrae population with a higher fraction of rare High Amplitude Short Period pulsators (see Fiorentino et al., 2015; Belokurov et al., 2018a). Given its properties, this centrally concentrated RR Lyrae population is likely of in-situ origin.
@@ 3.3 Chemical trends
Even though the GS/E progenitor is no longer intact, its pre-disruption properties can be studied now, provided the constituent stars are identified (see Sections 3.1 and  3.2). As the GS/E stars flood the Solar neighourhood, as aptly pointed out by Hasselquist et al. (2021), the community had been exploring the GS/E’s chemical trends long before we had realised these stars were all of common origin (see Nissen and Schuster, 2010; Hawkins et al., 2015; Fernández-Alvar et al., 2018; Hayes et al., 2018). Thus, the main trends, i.e. the downward slope of the $\alpha$-[Fe/H] track (Nissen and Schuster, 2010) and the constant low [Al/Fe] ratio (Hawkins et al., 2015) were known before the Gaia data started to arrive. These trends are updated, improved upon, and contrasted with those in surviving dwarf galaxies in Hasselquist et al. (2021) (see Fig. 13). They show that compared to the four most massive current satellites of the Milky Way, namely the LMC, the SMC, Sgr (holding on for dear life), and Fornax, the GS/E progenitor attains demonstrably higher [Mg/Fe] at $-1.5<$[Fe/H]$<-1$ (but still some 0.1 dex below the disc sequence at the same metallicity). Clearly, the GS/E dwarf was more efficient at converting gas into stars compared to the present-day survivors. Hasselquist et al. (2021) hypothesize that this is due to its proximity to the Milky Way. Equally, the environment may have also played a crucial role in shaping the SFHs of the Magellanic Clouds which appear to have lived a mostly dull and lethargic life in the suburbs of the Galaxy. However, having formed next to a bigger galaxy did not end well for the GS/E progenitor: its life was cut short, although none of the models in the literature reveal signs of abrupt truncation in star formation (possibly due to the rigidity of the SFH approximation).
The chemical trends of the GS/E can also be compared to other known debris in the stellar halo. For example, Horta et al. (2023b) compare the chemical abundances of several known substructures and find that some (namely, Arjuna, LMS-1, I’itoi, and perhaps Sequoia — see also Monty et al. 2020; Feuillet et al. 2021) are chemically indistinguishable from the GS/E, perhaps suggesting that they share a common origin. Thus, similarly to the dynamical association of the VOD and HAC with the GS/E (see Section 3.2), care must be taken in defining distinct substructures in the halo. Indeed, a massive structure like the GS/E is omnipresent, and can have several sub-components.
@@
Curiously, the first constraints on the details of the GS/E SFH based on the behaviour of its stars in the $\alpha$-[Fe/H] plane are also pre-Gaia (Fernández-Alvar et al., 2018). This can be compared to the models of Vincenzo et al. (2019) and Sanders et al. (2021) that were built to reproduce a much more fine-tuned selection of the likely GS/E members. All three models agree on the relatively short SFH of the GS/E progenitor: star formation activity had to stop after 3-4 Gyr because by then the dwarf was stripped of its gas and fully disassembled by the Milky Way tides. The star formation efficiency of Fernández-Alvar et al. (2018) and Vincenzo et al. (2019) are similar at 0.3-0.4 Gyr-1, whilst that in Sanders et al. (2021) is almost an order of magnitude lower at $\approx 0.06$ Gyr-1. To infer the SFH of the dwarf, the former two studies use theoretical elemental yields, while the latter simultaneously and self-consistently models a wide range of elements and pins down both the yields and the global parameters of the chemical enrichment model. The amplitude of the resulting star formation rate in the model of Vincenzo et al. (2019) is 3 $M_{\odot}$yr-1, some 7 times that of the model in Sanders et al. (2021). Integrating the SFR over GS/E’s lifetime, Vincenzo et al. (2019) find a total stellar mass of $\approx 5\times 10^{9}M_{\odot}$. Although not quoted directly in Sanders et al. (2021), their model would infer an order of magnitude lower total stellar mass of $\approx 5\times 10^{8}M_{\odot}$. Hasselquist et al. (2021) provide two models of the GS/E chemical evolution, one with a very broad SFH lasting more than 10 Gyr and one with a much narrower peak in agreement with the works above. Even though trained only on the [Mg/Fe]-[FeH] data, this latter model has a peak SFR much closer to the results of Sanders et al. (2021).
@@
Most of the chemical studies above use the exquisite APOGEE data. However, APOGEE is somewhat limited in its ability to measure heavy element abundances, and, in particular, has only scarce data on neutron-capture tracers (only Ce is listed amongst the relatively reliably measured species as part of the APOGEE DR17). The information on $s$- and $r$-process had to come from elsewhere, namely from the data releases of the GALAH survey (De Silva et al., 2015; Buder et al., 2018, 2021) and individual follow-up studies. These challenging measurements (due to a small number of weak lines) yielded a result not many had predicted: a noticeable enhancement in the $r$-process abundances in the GS/E stars compared to the rest of the halo (including the in-situ stars where available at comparable metallicities) and most of the surviving dwarf satellites (Aguado et al., 2021; Matsuno et al., 2021). Measuring the rather low [Ba/Eu]$\approx 0$ in their follow-up VLT spectra, Aguado et al. (2021) deduce that the neutron-capture production in the GS/E progenitor proceeded mostly via the $r$-process channel. Note that while [Eu/Fe]$\approx 0.65$ and [Eu/Mg]$\approx 0.3$ are at the upper end of what is observed, these values are still noticeably below the extreme $r$-process enhancement detected in one ultra-faint dwarf, namely Reticulum 2 (Ji et al., 2016) which reaches [Eu/Fe]$>1$, but similar to levels in other $r$-process enhanced UFDs such as Tucana 3 (Hansen et al., 2017) and Grus 2 (Hansen et al., 2020). Focusing on the tracks in [Eu/Mg]-[Mg/Fe] space for the GS/E stars with [Fe/H]$>-1.5$ in GALAH, Matsuno et al. (2021) conclude that the bulk of $r$-process in the dwarf must be due to neutron star mergers. Naidu et al. (2022) observe GS/E stars together with targets they designate “Kraken” with Magellan and measure (amongst other lines) Eu, Ba, and Mg abundances. At similar [Mg/H], the GS/E has 0.3 dex higher [Eu/Mg]. Naidu et al. (2022) assume that the “Kraken” had a more rapid star formation activity and conclude that the enrichment due to the neutron star mergers must be delayed by some 0.5 Gyr. Note that most “Kraken” stars analysed in this study would also pass the cuts for the Milky Way pre-disc in-situ population (Aurora — see 4.3.1). Not only do these stars have low orbital energy and high Mg abundance, 16 out of 20 objects in the sample lie above the accreted sequence in the plane of [Al/Fe] vs [Fe/H] (see panels 2 and 3 in Fig. 2 of Belokurov and Kravtsov 2022) and along the rapidly rising in-situ track (see Fig. 2 in Horta et al. (2021). In our opinion, replacing “Kraken” with “early Milky Way”, the arguments presented in Naidu et al. (2022) remain valid, or perhaps, become even more coherent.